Geometry. A square's side is increased by 9 yards, which corresponds to an increase in the area by 261 square yards. How many yards is the side of the initial square?
The sides of the square were ______ yards to begin with.

Answer by patrickca(4)

Area of square = s * s or s^2
Area of rectangle after increase = (s + 9)(s + 9) or s^2 + 18s + 81
SO...
s^2 + 261 = s^2 + 18s + 81
261 = 18s + 81
180 = 18s
10 = s